OpenStudy (mathlegend):

Solve for p: A = 2πpw

OpenStudy (ghazi):

divide both sides by @ *pi*w

OpenStudy (mathlegend):

πw? why not the 2π

OpenStudy (ghazi):

\[\frac{ A }{ 2 \pi w }=\frac{ 2 \pi w p }{ 2 \pi w p }=p\]

OpenStudy (ghazi):

sorry it is 2, pi and w

OpenStudy (ghazi):

@MathLegend clear?

OpenStudy (mathlegend):

Yes, thanks @ghazi

OpenStudy (ghazi):

:)